OpenStudy (anonymous):

Simplify the expression below: Log((10^6)(X^2))

OpenStudy (anonymous):

A 6+2log(x) B 6x^2 C 12 D 12log(x)

Nnesha (nnesha):

familiar with the log rules ?

OpenStudy (anonymous):

Not at all

OpenStudy (anonymous):

Any type of mathematics are not my best subjects.

Nnesha (nnesha):

quotient rule\[\large\rm log_b x - \log_b y = \log_b \frac{ x }{ y}\] to expand you can change division to subtraction product rule \[\large\rm log_b x + \log_b y = \log_b( x \times y )\] addition < ----> multiplication power rule \[\large\rm log_b x^y = y \log_b x\] that's all you need to know for this equation these are log rules

Nnesha (nnesha):

this topic is very easy they gave you 2 rules you just need to follow em

Nnesha (nnesha):

\[\huge\rm Log( 10^6 X^2 )\]which rule would you apply ??

Nnesha (nnesha):

\[\huge\rm Log( 10^6* X^2 )\]which rule would you apply ?? multiplication so change that to ??

OpenStudy (anonymous):

I dont know, this is like a foreign language to me, Im sorry.

Nnesha (nnesha):

look at the rules i gave you try it !

Nnesha (nnesha):

here is an example \[\log (2 y) \] its `2 times y` multiplication so to expand i would apply product rule \[\rm \log(2)+\log(y)\]split it into two log

OpenStudy (anonymous):

Log(10) and log(x)? or would the exponents come before the Log, 6Log(10) and 2Log(x) like this?

Nnesha (nnesha):

yes that's correct !

Nnesha (nnesha):

but wait what sign would be between both logs ?

Nnesha (nnesha):

product rule \[\large\rm log_b(x\color{reD}{*}y)= log_b x\color{ReD}{ +} \log_b y \] addition < ----> multiplication to expand change multiplication to addition

OpenStudy (anonymous):

okay, so Log(10)+Log(x) is what it looks like then?

OpenStudy (anonymous):

this is confusing, sorry about this.

Nnesha (nnesha):

yes don't forget the exponents \[\rm Log(10)^6+Log(x) ^2\] now apply power rule move the exponents front of the logs

Nnesha (nnesha):

it's okay this is the easy subject :=))

Nnesha (nnesha):

i mean topic*

OpenStudy (anonymous):

okay so 6Log(10)+2Log(x).

Nnesha (nnesha):

loks good!

Nnesha (nnesha):

looks*

OpenStudy (anonymous):

Thank you so much, this helped alot.

Nnesha (nnesha):

np:=))